• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name milbrath

    Log entry time 00:04:34 on May28,2000

    Entry number 42130

    keyword=shift summary

    16:00-24:00 May-27-00
    Milbrath, Breuer, Sirca

    Target was restored at end of previous shift. A harp scan was performed at the start of shift.
    Target recovery work:
    Singles studies electron angle = 18.067 deg, hadron angle = 34.286 deg, hadron P0 = 1.552 16:13 - 16:50 runs 1346-9
    Singles studies hadron angle to 47.089 deg 17:10 - 17:35 runs 1350-2
    Coins 17:36 - 18:54 runs 1353-7
    Carbon singles 18:55-19:05 run 1358
    Singles studies electron angle to 14.074 deg 19:10 - 19:40 runs 1359-63
    Access to reboot hadron vernier camera power supply 19:45-20:15
    harp scan 20:15 - 20:20 run 1364
    retuning of beam 20:20-21:00
    Single studies hadron angle to 34.29, hadron Po to 1.270 21:00- 21:55
    runs 1365-71

    Production data for theta(cm) = 50 deg, phi = 0 deg 22:00 - runs 1372 - 1376 (at midnight)

    DETAILS:
    Singles studies, electron = 18.067 deg, hadron = 34.286 deg, P0 = 1.552
    1346 - carbon
    1347 - 15cm dummy
    1348 - 10cm dummy
    1349 - 4cm dummy

    Singles studies, electron = 18.067 deg, hadron = 47.089 deg, P0 = 1.552:
    1350 - 4cm dummy
    1351 - 15cm dummy
    1352 - 10cm dummy

    Coins, same as above:
    1353,4 - LH2 coins only - 6 msr collimators
    1355 - LH2 coins only -sieve collimator for electron, spot moved
    1356 - no good
    1357 - LH2 coins only - sieve collimator of electron

    1358 - carbon - singles - 6msr collimators back

    Singles studies, electron = 14.074 deg, hadron = 47.089 deg, P0 = 1.552:
    1359 -carbon singles - too high deadtime, short
    1360 - better prescale rate
    1361 - 4 cm dummy
    1362 - 15cm LH2 -short junk run
    1363 - 15cm LH2

    1364 - harp scan

    Singles studies, electron = 14.074 deg, hadron = 34.290 deg, P0 = 1.27 GeV/c:
    1365 - junk
    1366 - carbon target
    1367 - 4 cm dummy
    1368 - 15 cm dummy
    1369 - 10 cm dummy
    1370 - 15cm LH2 -prescales too high
    1371 - 15cm LH2

    Production runs (same kinematics as above):
    1372 - 1 million events
    1373 - 1.1 million events
    1374 - 1.1 million events
    1375 - 1.1 million events
    1376 -